James invested Php20,000 for one year and earned Php1470 interest. if part of the money is invested at 10%and the remainder is invested at 6%, how much is invested at each rate?

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52781)

.
James invested Php20,000 for one year and earned Php1470 interest.
if part of the money is invested at 10%and the remainder is invested at 6%,
how much is invested at each rate?
~~~~~~~~~~~~~~~~~~

x         dollars invested at 10%,
(20000-x) dollars invested at  6%.


The total annual interest equation is

    0.1x + 0.06(20000-x) = 1470  dollars.


Simplify and find x

    0.04x + 1200 = 1470

    0.04x = 1470 - 1200

    0.04x = 270

        x = 270/0.04 = 6750.


ANSWER.  $6750 invested at 10%;  the rest 20000 - 6750 = 13250 dollars invested at 6%.


CHECK.   0.1*6750 + 0.06*13250 = 1470 dollars, total annual interest.  ! correct !

Solved.

--------------------

To see many other similar solved problems using one equation in one unknown,  look into the lesson
    - Typical investment problems
in this site.

To see many other similar solved problems using systems of two equations in two unknowns,  look into the lesson
    - Using systems of equations to solve problems on investment
in this site.



Answer by greenestamps(13200)



The response from the other tutor shows a solution using the standard formal algebraic method.

Here is an informal method that can be used to solve any 2-part "mixture" problem like this. This method is fast and easy, especially if the numbers are "nice".

Php1470 interest on Php20,000 is an interest rate of 1470/20000 = 0.0735 = 7.35%.

The ratio in which the total was split between the two investments is exactly determined by where that 7.35% interest rate lies between the individual interest rates of 6% and 10%.

So look at the three percentages 6, 7.35, and 10 (on a number line, if it helps), and observe/calculate that the 7.35% interest rate is 1.35/4 = 135/400 = 27/80 of the way from 6% to 10%.

That means 27/80 of the total was invested at the higher rate.

20000*(27/80) = 27*(20000/80) = 27*250 = 6750

ANSWERS: Php6750 was invested at 10%; the other Php13250 was invested at 6%

CHECK: .10(6750)+.60(13250) = 675+795 = 1470